Dont tell me weve lost another bid! exclaimed Janice Hudson, president of Prime Products Inc. Im afraid so, replied Do2. Suppose that instead of using a plantwide overhead rate, the company had used a separate predetermined overhead rate in ea5. At the end of the year, the company assembled the following actual cost data relating to all jobs worked on during the yea
Answer #1

Part 1 –a

Predetermined MOH rate = estimated total MOH / estimated total amount of the allocation base = 1490400/931500 = 160% of direct labor cost

Part 1-b

Applied MOH = Predetermined MOH rate * actual use of the allocation base to determine rate = 160%*30300 = $48480

Part 2-a

Predetermined MOH rate = estimated total MOH / estimated total amount of the allocation base

MOH rate for Cutting

558900/310500

180% of direct labor cost

MOH rate for Machining

828000/207000

400% of direct labor cost

MOH rate for Assembly

103500/414000

25% of direct labor cost

Part 2-b

Applied MOH = Predetermined MOH rate * actual use of the allocation base to determine rate

Applied MOH Cutting

10000*180%

$18000

Applied MOH Machining

3100*400%

$12400

Applied MOH Assembly

17200*25%

$4300

Total MOH applied

$34700

                                               

Part 4-a

Direct materials

26200

Direct labor

30300

Manufacturing overhead

48480

Total manufacturing costs

$104980

Markup

150%

Bid price

$157470

Part 4-b

Direct materials

26200

Direct labor

30300

Manufacturing overhead

34700

Total manufacturing costs

$91200

Markup

150%

Bid price

$136800

Part 5-a

Applied overhead = 905700*160% = $1449120

Underapplied overhead = actual overhead – applied overhead = 1551000-1449120 = $101880

Part 5-b

Applied overhead = (334000*180%)+(220500*400%)+(351200*25%) = $1571000

Overapplied overhead = applied overhead – actual overhead = 1571000-1551000 = $20000